OpenStudy (anonymous):

Solve this problem without using a similar triangles approach: A street light is mounted at the top of a 15 ft tall pole. A man 6 ft tall walks away from the pole with a speed of 5 ft/s along a straight path. How fast is the tip of his shadow moving when he is 40 ft from the pole?

hero (hero):

Draw it. There's drawing function below

hero (hero):

If you can't use similar triangles, what approach are you supposed to use?

OpenStudy (anonymous):

is there a way to find it without using it?

hero (hero):

Not really sure. Why?

OpenStudy (anonymous):

 well that's what im trying to find. i understand it using similar triangles, but im trying to find another approach.

hero (hero):

Why would you want to use another approach? What's wrong with the one you already have?

OpenStudy (anonymous):

i tried breaking it up using the area of a triangle and adding it to the area of a trapazoid, to the find the area of total. however, when i took the derivative, i had too many missing variables, including dy/dt. that's the question im being asked.

OpenStudy (anonymous):

\[A _{total} = h ((b _{1} + b _{2}) / 2) + (1/2 * bh)\] i tried using this. when i plugged what i had into the derivative of this, it looked like this: \[d/d _{t} = 21/2 *5 + 3*d _{y}/d _{t}\] dx/dt was 5ft/s, the given speed of the man. does this work?

hero (hero):

Ask Satellite

hero (hero):

Okay, I can get you an answer but it will take a little time

hero (hero):

But I can't promise that the solution will not include similar triangles

OpenStudy (anonymous):

oh ok. thanks. instead of using that equation i gave above, i just used the velocity formula knowing that it's the derivative of an equation. \[V = x / t\] where it becomes in terms of d/dt: \[dV/dt = dx/dt \div d/dt\] substitute what is given, and i got: \[d/dt = 40/5\] \[d/dt = 8ft/s\] the answer in my book is 25/3 ft/s, which equals roughly 8.3 ft/s